Find the monthly payment and estimate the remaining balance. Assume interest is on the unpaid balance. Twenty​-year mortgage for ​$220,000 at 4.02​%; remaining balance after 12 years. The monthly payment is ​$.............. ​(Round to the nearest cent as​ needed.)
Answer #1

(a) Monthly Payment:

Loan amount, L =   $        220,000.00

Interest Rate, R % = 4.02

Remitting Frequency per year, N = 12

Time in Years, T = 22

Interest Rate per Period, ι = R/100N = 0.0034

Number of Periods, p = NT = 264

Installment, PMT = L * [ι + {ι / {((1 + ι)^p) - 1}}] = $1,256.76

(b) Amortization schedule:

After Monthly Payment Number Total paid Total paid towards Interest Total paid towards Principal Balance payable
1 $1,256.76 $737.00 $519.76 $219,480.24
2 $2,513.53 $1,472.26 $1,041.27 $218,958.73
3 $3,770.29 $2,205.77 $1,564.52 $218,435.48
4 $5,027.05 $2,937.53 $2,089.52 $217,910.48
5 $6,283.81 $3,667.53 $2,616.28 $217,383.72
6 $7,540.58 $4,395.76 $3,144.81 $216,855.19
7 $8,797.34 $5,122.23 $3,675.11 $216,324.89
8 $10,054.10 $5,846.92 $4,207.18 $215,792.82
9 $11,310.87 $6,569.82 $4,741.04 $215,258.96
10 $12,567.63 $7,290.94 $5,276.69 $214,723.31
11 $13,824.39 $8,010.26 $5,814.13 $214,185.87
12 $15,081.15 $8,727.79 $6,353.37 $213,646.63
13 $16,337.92 $9,443.50 $6,894.41 $213,105.59
14 $17,594.68 $10,157.41 $7,437.27 $212,562.73

...

140 $175,946.79 $83,298.28 $92,648.52 $127,351.48
141 $177,203.56 $83,724.90 $93,478.65 $126,521.35
142 $178,460.32 $84,148.75 $94,311.57 $125,688.43
143 $179,717.08 $84,569.81 $95,147.27 $124,852.73
144 $180,973.84 $84,988.06 $95,985.78 $124,014.22

Remaining balance = $124,014.22